OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. -20, -16, -12, -8, ...

OpenStudy (cwrw238):

nth term = a1 + (n-1)d where a1 = first term and d = common difference what is d for this sequence?

OpenStudy (anonymous):

d=4 -20 + (n-1)4

OpenStudy (anonymous):

Sp for this one it would be: Find an equation for the nth term of the arithmetic sequence. -1, 2, 5, 8, an= -1+(n-1)3 ?

OpenStudy (cwrw238):

yes